OpenStudy (anonymous):

In the function f(x) = 3x + 9 how does a decrease of 1 in the independent variable x affect f(x)? f(x) will have a decrease of 3 f(x) will have a decrease of 12 f(x) will have an increase of 3x + 3 f(x) will have an increase of 3

OpenStudy (radar):

f(x-1)=3(x-1) +9 3x-3+9 3x+6 f(x)=3x+9 decrease = 3x+6 ------- 3

OpenStudy (radar):

You better get a grip on these if you intend to complete calculus.